13N decays by positron emission. Balance the following nuclear equation by giving the mass number, atomic number, and element symbol for the missing species.
Answer #1

Skeleton equation: \({ }_{?}^{\mathrm{1}} \mathrm{N} \longrightarrow{ }_{z}^{A} X+{ }_{+1}^{0} e\)

Sum of the atomic numbers on both sides of the equation should be equal.

$$ \begin{aligned} &7=Z+1 \\ &Z=6 \end{aligned} $$

Thus, the atomic number of the new niclide is \(6 .\)

According to the periodic table, el ement with the atomic number 6 is carbon \((C)\).

Sum of the mass numbers on both sides of the equation should be equal.

$$ \begin{gathered} 13=A+0 \\ A=13 \end{gathered} $$

Thus, the mass number of the new niclide is 13 .

Therefore, the missing nuclide is \({ }_{6}^{13} \mathrm{C}\) Balanced equation: \({ }_{7}^{\mathrm{B}} \mathrm{N} \longrightarrow{ }_{6}^{13} \mathrm{C}+{ }_{+1}^{0} e\)